OpenStudy (anonymous):

Expand the series and evaluate: ∑ {5(on top), k=1(on bottom), k^3 on the right side}

OpenStudy (anonymous):

\[\sum_{k=1}^{5} k^3\]

OpenStudy (amistre64):

i think its just the square of k^1

OpenStudy (amistre64):

\[\sum k=\frac{n(n+1)}{2}\] \[\sum k^3=\left(\frac{n(n+1)}{2}\right)^2\] rings a bell to me

OpenStudy (across):

It's always a good idea to list out the elements of series this small:\[1^3+2^3+3^3+4^3+5^3.\]

OpenStudy (across):

Just to get a feel for it, you feel me!? ^^

OpenStudy (anonymous):

Yeah, thanks. So would it be 1, 8, 27, 64, 125?

OpenStudy (across):

Don't forget that it's a summation! :P

OpenStudy (anonymous):

What do you mean?

OpenStudy (turingtest):

you have to add them...

OpenStudy (anonymous):

Deriving this proof is a good excercise, there are many ways actually, but I like way of differentiating and then substituting .. probably that's how euler did it!

OpenStudy (amistre64):

euler plagarized the chinese :P

OpenStudy (anonymous):

^^ lol, how do you know ? :P

OpenStudy (amistre64):

past life regression :)

OpenStudy (amistre64):

if we take the derivative of this life ....

OpenStudy (anonymous):

You were there when Euler did it? ... omG! :D

OpenStudy (amistre64):

yep, just call me JC Superstar yay!!

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

Okay, I hope you two know that that wasn't very helpful. I am just wondering how to solve this. Across, what do you mean by a summation?

OpenStudy (turingtest):

ADD them up! that is what summation means

OpenStudy (turingtest):

I said it above but nobody cared...

OpenStudy (turingtest):

\[1^3+2^3+3^3+4^3+5^3=?\]